### v1

# Skew Laurent series and general cyclic convolutional codes

Abstract.
Convolutional codes were originally conceived as vector subspaces of a finite-dimensional vector space over a field of Laurent series having a polynomial basis. Piret and Roos modeled cyclic structures on them by adding a module structure over a finite-dimensional algebra skewed by an algebra automorphism. These cyclic convolutional codes turn out to be equivalent to some right ideals of a skew polynomial ring built from the automorphism. When a skew derivation is considered, serious difficulties arise in defining such a skewed module structure on Laurent series. We discuss some solutions to this problem which involve a purely algebraic treatment of the left skew Laurent series built from a left skew derivation of a general coefficient ring, when possible.
Research funded by the IMAG–María de Maeztu grant (CEX2020-001105-M) and “Proyecto PID2023-149565NB-100” from MICIU/AEI /10.13039/501100011033 and FEDER, UE”
Introduction
Let $\mathbb{F}(\!(X)\!)$ be the Laurent formal power series field with coefficients in a field $\mathbb{F}$. Classically, see e.g. [8], $\mathbb{F}$–linear convolutional codes of length $n$ may be conceived as $\mathbb{F}(\!(X)\!)$–vector subspaces $\mathcal{D}$ of $\mathbb{F}^{n}(\!(X)\!)$ generated by vectors with entries in the rational field $\mathbb{F}(X)$ or, equivalently, with an $\mathbb{F}^{n}(\!(X)\!)$–basis of elements of $\mathbb{F}^{n}[X]$. In this framework, the indeterminate $X$ is interpreted as the shift operator acting on blocs $v_{i}\in\mathbb{F}^{n}$ as $Xv_{i}=v_{i-1}$. It turns out that the map that sends $\mathcal{D}$ to $\mathcal{D}\cap\mathbb{F}^{n}[X]$ is a bijection between convolutional codes and $\mathbb{F}[X]$–submodules of the free $\mathbb{F}[X]$–module $\mathbb{F}^{n}[X]$ that are direct summands.
Cyclic structures on convolutional codes were initially introduced by adding a module structure on $\mathbb{F}^{n}(\!(X)\!)$. More concretely, identifying vectors of $\mathbb{F}^{n}$ with polynomials in $\mathbb{F}[t]$ of degree up to $n-1$, and taking the quotient ring $A=\mathbb{F}[t]/\langle t^{n}-1\rangle$, then a right [^1] In [8], a left $A$–module structure is considered. We prefer to twist left and right hands in our approach since our references on skew derivations fit better this equivalent formalism. $A$–module structure on $\mathbb{F}^{n}(\!(X)\!)$ is defined. To avoid trivial cases, as noted in [7], such a module structure is twisted by an $\mathbb{F}$–algebra automorphism $\sigma$ of $A$. Then, a convolutional code is cyclic if and only if it is a right $A$–submodule of $\mathbb{F}^{n}(\!(X)\!)$ ([8, Theorem 3]).
Under the aforementioned bijection, cyclic convolutional codes correspond to right ideals of $A[X;\sigma]$ that are $\mathbb{F}[X]$–direct summands (see [4, Proposition 2] for a more general statement). Here, $A[X;\sigma]$ is the left skew polynomial ring defined from $\sigma$. Indeed, this is the way in which cyclic convolutional codes are defined in [3]. More generally, cyclic convolutional codes are conceived in [6] as right ideals of the left skew polynomial ring $A[X;\sigma,\delta]$ built from a left skew derivation $(\sigma,\delta)$ of any finite-dimensional $\mathbb{F}$–algebra $A$. When trying to return to Roos’ viewpoint, a major difficulty is that no obvious way to give a right $A$–module structure on $A(\!(X)\!)$, twisted by $(\sigma,\delta)$, is available. For $\delta=0$, giving such a structure, as in [8], is equivalent to endow $A(\!(X)\!)$ of a ring structure extending that of $A[X;\sigma]$, which is always possible if $\sigma$ is an automorphism, namely, the left skew Laurent power series ring $A(\!(X;\sigma)\!)$. When $\delta\neq 0$, serious difficulties arise, being a traditional solution in the realm of Ring Theory, to consider inverse skew Laurent formal power series ring (see, e.g. [2, Section 2.3]). However, this object does not fit in the described picture of cyclic convolutional codes.
Although the ring of left skew Laurent formal power series $A(\!(X;\sigma,\delta)\!)$ could not, with the desired properties, exist for a general $\delta\neq 0$, suitable restrictions on the left skew derivation $(\sigma,\delta)$ allows to define a well founded product extending that of $A[X;\sigma,\delta]$. Thus, in [10], under the hypotheses that $\sigma$ is an automorphism and both $\delta,-\delta\sigma^{-1}$ are nilpotent, besides additional topological conditions on the ring $A$ and $(\sigma,\delta)$, such a ring structure is modeled. On the other side, under the condition $\delta\sigma=q\sigma\delta$ for some $q\in\mathbb{F}$ for $A$ an $\mathbb{F}$–algebra, the ring of left skew formal power series $A[\![X;\sigma,\delta]\!]$ is defined in [1] whenever $\delta$ is locally nilpotent. For a general coefficient ring $A$, and with no condition apart from being locally nilpotent, a well defined product for formal power series with coefficients in $A$ is obtained in [5].
In [4], for $\sigma$ an algebra automorphism of a finite-dimensional $\mathbb{F}$–algebra $A$ and $\mathbb{F}^{n}$ endowed with the structure of a right $A$–module, $(\sigma,A)$–cyclic convolutional codes are defined, in the spirit and generalizing Roos’ notion, as right $A$–submodules of $\mathbb{F}^{n}(\!(X)\!)$. The bijective correspondence between these $(\sigma,A)$–cyclic convolutional codes and right $A[X;\sigma,\delta]$–submodules of $\mathbb{F}^{n}[X]$ that are $\mathbb{F}[X]$–direct summands is proved in [4, Proposition 2]. With the aim of exploring to what extent this correspondence still is possible for a left skew derivation $(\sigma,\delta)$ on $A$, we analyzed how the skew product defined in [5] could be extended to Laurent series. Although we are mainly interested, in the realm of convolutional codes, in the case of a finite-dimensional algebra $A$, our analysis is done on a left skew derivation on a general ring $A$. We think the outcome could be of independent interest, and it is explained in Section 2.
In order to extend the notion of $(\sigma,A)$–cyclic convolutional code to that of $(\sigma,\delta,A)$–cyclic convolutional code, when possible, Section 3 collects the needed technical facts on modules over $A[\![X;\sigma,\delta]\!]$ and $A(\!(X;\sigma,\delta)\!)$.
In Section 4 we define, starting from a right $A$–module structure on $\mathbb{F}^{n}$, when possible, a general concept of cyclic convolutional code modeled by a left skew derivation $(\sigma,\delta)$ on a finite-dimensional $\mathbb{F}$–algebra $A$. We call these convolutional codes $(\sigma,\delta,A)$–cyclic and, when $\delta=0$, they boil down to the $(\sigma,A)$–cyclic convolutional codes from [4] and, ultimately, to cyclic convolutional codes in the sense of Roos. The bijective correspondence between $(\sigma,\delta,A)$–cyclic convolutional codes and $\mathbb{F}[X]$–direct summands that are right $A[X;\sigma,\delta]$–submodules of $\mathbb{F}^{n}[X]$ is proved in Theorem 4.8. As a consequence, we deduce that the notions of $(\sigma,\delta,A)$–cyclic convolutional code and (right) ideal code from [6] are equivalent, whenever $\delta$ and $\delta^{\prime}$ are nilpotent.
Some examples are provided along the paper.
1. Notation for formal power series
Let $M$ be a left module over a ring $A$ and consider the left $A$–module $M^{\mathbb{Z}}$ of all maps from $\mathbb{Z}$ to $M$. We will use the sequence-like notation $(u_{i})_{i\in\mathbb{Z}}$ for these maps.
In the realm of convolutional codes, representing these sequences as formal power series
$\sum_{i\in\mathbb{Z}}u_{i}X^{i}=\sum_{i=-\infty}^{\infty}u_{i}X^{i},$
in an indeterminate $X$ turns out to be useful. Indeed, the delay operator $X$ on $M^{\mathbb{Z}}$ is defined by the rule $X(u_{i})=u_{i-1}$. This is obviously an $A$–linear invertible operator.
Under this formalism, the left $A$–module $M^{\mathbb{Z}}$ will be denoted by $M(\!(X,X^{-1})\!)$.
Some of its relevant submodules are the formal Laurent series
$M(\!(X)\!)=\left\{\sum_{i=i_{0}}^{\infty}u_{i}X^{i}:i_{0}\in\mathbb{Z}\right\},$
the Laurent polynomials
$M[X,X^{-1}]=\left\{\sum_{i=i_{0}}^{i_{1}}u_{i}X^{i}:i_{0},i_{1}\in\mathbb{Z},i_{0}\leq i_{1}\right\},$
and the polynomials
$M[X]=\left\{\sum_{i=0}^{n}u_{i}X^{i}:n\in\mathbb{N}\right\},$
all of them with coefficients in $M$. We will also consider the $A$–submodules of all formal power series
$M[\![X]\!]=\left\{\sum_{i=0}^{\infty}u_{i}X^{i}\right\}.$
With $A$ a commutative ring and $M=A$, all these $A$–modules, with the exception of $A(\!(X,X^{-1})\!)$, become commutative rings in the usual way. Even in this case, there are other possible structures of ring on these additive groups. So, in general, we do not assume a priori any ring structure on them.
2. Skew Laurent formal power series
A left skew derivation of a ring $A$ is a pair $(\sigma,\delta)$ of additive maps such that $\sigma$ is a ring endomorphism of $A$ and $\delta$ satisfies the equality $\delta(ab)=\sigma(a)\delta(b)+\delta(a)b$ for all $a,b\in A$. A right skew derivation over $A$ is just a left skew derivation over the opposite ring $A^{{}^{\prime}}$ of $A$.
Given a left skew derivation $(\sigma,\delta)$ over $A$, the free left $A$–module $A[X]$ with basis $\{X^{i}:i\in\mathbb{N}\}$ is endowed with a ring structure whose product is determined by the rules
(2.1) $X^{i+j}=X^{i}X^{j},\quad Xa=\sigma(a)X+\delta(a),\qquad(i,j\in\mathbb{N},a\in A).$
This ring is denoted by $A[X;\sigma,\delta]$ and their elements are called (left) skew polynomials, and it is referred to as the skew polynomial ring of $(\sigma,\delta)$. A symmetric construction applies for any right skew derivation $(\sigma^{\prime},\delta^{\prime})$ of $A$. We will use the notation $[X;\sigma^{\prime},\delta^{\prime}]A$ in this case.
Given any additive map $L:A\to A$ and $f=\sum_{i}f_{i}X^{i}\in A[X;\sigma,\delta]$, set $L(f)=\sum_{i}L(f_{i})X^{i}$. We easily get
(2.2) $Xf=\sigma(f)X+\delta(f).$
The multiplication rule (2.2) extends to
(2.3) $X^{n}f=\sum_{k=0}^{n}N_{k}^{n}(f)X^{k},$
for all $f\in A[X;\sigma,\delta]$, for additive endomorphisms of $A[X;\sigma,\delta]$ defined recursively, for $0\leq i\leq n$, by
$N_{i}^{n+1}=\sigma N_{i-1}^{n}+\delta N_{i}^{n},\qquad(N_{-1}^{n}=0,N_{0}^{0}=id_{A[X;\sigma,\delta]},N_{n+1}^{n}=0).$
It follows from (2.3) that
(2.4) $\sum_{i=0}^{n}X^{i}a_{i}=\sum_{i=0}^{n}\left(\sum_{j=i}^{n}N_{i}^{j}(a_{j})\right)X^{i},$
for all $a_{0}\dots,a_{n}\in A$, and
(2.5) $\left(\sum_{i=0}^{n}g_{i}X^{i}\right)f=\sum_{i=0}^{n}\left(\sum_{k=i}^{n}g_{k}N_{i}^{k}(f)\right)X^{i},$
for all $g_{0},\dots,g_{n}\in A,f\in A[X;\sigma,\delta]$.
When $\sigma$ is an automorphism, $(\sigma^{-1},-\delta\sigma^{-1})$ is a right skew derivation, which will be denoted by $(\sigma^{\prime},\delta^{\prime})$. Observe that, for all $f\in A[X;\sigma,\delta]$, one deduces from (2.2) that
(2.6) $fX=X\sigma^{\prime}(f)+\delta^{\prime}(f),$
for all $f\in A[X;\sigma,\delta]$. We get from (2.4) and (2.6) that $\{X^{i}:i\in\mathbb{N}\}$ is a basis of $A[X;\sigma,\delta]$ as right $A$–module, which implies, by (2.6), that $A[X;\sigma,\delta]=[X;\sigma^{\prime},\delta^{\prime}]A$.
A definition of a ring structure on $A[\![X]\!]$ extending that of $A[X;\sigma,\delta]$ might not be possible. We are interested in reasonable products according to Definition 2.1 below. Set, for any formal power series $s=\sum_{i=0}^{\infty}s_{i}X^{i}$ with coefficients $s_{i}\in A$, and any additive endomorphism $L:A\to A$,
$L(s)=\sum_{i=0}^{\infty}L(s_{i})X^{i}.$
Definition 2.1.
We will say that ring structure on $A[\![X]\!]$ is reasonable if $A$ is a subring of $A[\![X]\!]$, the left $A$–module of $A[\![X]\!]$ is given by restriction of scalars of the ring extension $A\leq A[\![X]\!]$ and the product of $A[\![X]\!]$ satisfies, for any formal power series $s$, the equality
(2.7) $\left(\sum_{i=0}^{\infty}s_{i}X^{i}\right)X=\sum_{i=0}^{\infty}s_{i}X^{i+1}.$
Lemma 2.2.
Given a reasonable ring structure on $A[\![X]\!]$ and $(\sigma,\delta)$ a left skew derivation of $A$, it turns out that $A[X;\sigma,\delta]$ becomes, with its usual product, a subring of $A[\![X]\!]$ if and only if
(2.8) $Xa=\sigma(a)X+\delta(a),$
for all $a\in A$.
Proof.
Obviously, $A[X]$ is a left $A$–submodule of $A[\![X]\!]$. If (2.8) holds then, $Xa\in A[X]$ for all $a\in A$. Inductively, by using (2.7), we get more generally that $XaX^{m}\in A[X]$ for every $m\geq 1$. We get thus that $Xg\in A[X]$ for any $g\in A[X]$. From this it is easily checked that $fg\in A[X]$ for every $f,g\in A[X]$. Therefore, $A[X]$ is a subring of $A[\![X]\!]$ which, by (2.8) is isomorphic to $A[X;\sigma,\delta]$. The converse is clear. ∎
Lemma 2.2 makes reasonable the following definition.
Definition 2.3.
Let $A[\![X]\!]$ be endowed with a reasonable product that satisfies (2.8). We then say that this ring, denoted by $A[\![X,\sigma,\delta]\!]$, is the ring of left skew formal power series with respect to $(\sigma,\delta)$. Shortly, we say just that the ring $A[\![X;\sigma,\delta]\!]$ exists.
It is not clear, in general, whether $A[\![X;\sigma,\delta]\!]$ exists. Under the assumption that $\sigma$ is an automorphism, in conjunction with suitable topological hypotheses on $A$, $\sigma$, and $\delta$, a reasonable ring structure is defined in [9]. When $\delta$ is assumed to be locally nilpotent, that is, for any $a\ \in A$, $\delta^{k}(a)=0$ for $k$ large enough, there is a well defined skew product of formal power series extending that of $A[X;\sigma,\delta]$, see [5, Theorem 5.3]). This last product requires no additional conditions. In the first case, the ring of skew Laurent formal power series, as a (left and right) ring of fractions of the skew power series ring is defined in [10], added that $\delta$ and $\delta^{\prime}$ are nilpotent.
Let us describe the product defined in [5], which subsumes that of [10] . According to [5, Lemma 5.2], if $\delta$ is locally nilpotent, then, for every $a\in A$ and $n\in\mathbb{N}$, there exists a natural number $O_{n}(a)$ such that, with the product of $A[X;\sigma,\delta]$,
(2.9) $(X^{m}a)_{i}=0\text{ whenever }i\leq n\text{ and }m\geq O_{n}(a).$
The skew product of $s,t\in A[\![X]\!]$ is then defined in [5, Theorem 5.3] as follows. For every $n\in\mathbb{N}$, set $q\geq\max_{0\leq i\leq n}\{O_{n}(t_{i})\}$ and
(2.10) $(st)_{n}=\left[\left(\sum_{i=0}^{q}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n},$
where the right hand product is done in $A[X;\sigma,\delta]$. This gives the $n$–th coefficient of the product series $st$, independently on the choice of $q$, by virtue of (2.9). A straightforward argument shows that this product on $A[\![X]\!]$ is reasonable and (2.8) holds.
Proposition 2.4.
If $\delta$ is locally nilpotent, then there exists a unique reasonable ring structure on $A[\![X]\!]$ containing $A[X;\sigma,\delta]$ as a subring. The product obeys (2.10) and, moreover,
(2.11) $Xs=\sigma(s)X+\delta(s),$
for every $s\in A[\![X;\sigma,\delta]\!]$.
Proof.
By [5, Theorem 5.3], the product on $A[\![X]\!]$ given by (2.10) is well defined. All conditions required to endow $A[\![X]\!]$ with the structure of a ring are clearly satisfied by this product, with the possible exception of the associativity. Let us check it. To this end, given $s\in A[\![X]\!]$ and $n\in\mathbb{N}$, set
$tr_{n}(s)=\sum_{i=0}^{n}s_{i}X^{i}\in A[X;\sigma,\delta],$
and
$q_{n}(s)=\max_{0\leq i\leq m\leq n}\{O_{m}(s_{i})\}.$
Given $s,t\in A[\![X]\!]$, we get from (2.9) that, for any $q\geq q_{n}(t)$
$tr_{n}(st)=\sum_{m=0}^{n}(st)_{m}X^{m}=\sum_{m=0}^{n}\left[\left(\sum_{i=0}^{q}s_{i}X^{i}\right)\left(\sum_{i=0}^{m}t_{i}X^{i}\right)\right]_{m}X^{m}$
$=\sum_{m=0}^{n}\left[\left(\sum_{i=0}^{q}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{m}X^{m},$
since
$\left[\left(\sum_{i=0}^{q}s_{i}X^{i}\right)\left(\sum_{i=m+1}^{n}t_{i}X^{i}\right)\right]_{m}=0.$
Therefore, for $q\geq q_{n}(t)$,
(2.12) $tr_{n}(st)=tr_{n}(tr_{q}(s)tr_{n}(t)).$
Finally, if $s,t,u\in A[\![X]\!]$, then, for $q\geq q_{n}(st)$, we have
$tr_{n}(u(st))=tr_{n}(tr_{q}(u)tr_{n}(st))=tr_{n}(tr_{q}(u)tr_{q_{n}(t)}(s)tr_{n}(t)),$
while, for $q\geq q_{q_{n}(t)}(s)$, we get
$tr_{n}((us)t)=tr_{n}(tr_{q_{n}(t)}(us)tr_{n}(t))=tr_{n}(tr_{q}(u)tr_{q_{n}(t)}(s)tr_{n}(t)).$
In this way, setting $q\geq q_{n}(st),q_{q_{n}(t)}(s)$, we see that $tr_{n}(u(st))=tr_{n}((us)t)$ for every $n$, which certainly implies that $u(st)=(us)t$, as desired.
This product clearly satisfies (2.7). If $\cdot$ is the product of some ring structure on $A[\![X]\!]$ extending that of $A[X;\sigma,\delta]$ and satisfying $(\sum_{i=0}^{\infty}s_{i}X^{i})\cdot X=\sum_{i=0}^{\infty}s_{i}X^{i+1}$ for every series $s$, then, clearly, $(s\cdot X^{m})_{n}=0$ for every $m>n\geq 0$. Given $s,t\in A[\![X]\!]$, and $n\in\mathbb{N}$, write
$s=\sum_{i=0}^{q_{n}(t)}s_{i}X^{i}+\sum_{i>q_{n}(t)}s_{i}X^{i},\quad t=\sum_{i=0}^{n}t_{i}X^{i}+\sum_{i>n}t_{i}X^{i}.$
Thus,
$(s\cdot t)_{n}=\left(s\cdot\sum_{i=0}^{n}s_{i}X^{i}\right)_{n}+\left(s\cdot\sum_{i>n}t_{i}X^{i}\right)_{n}\\ =\left(s\cdot\sum_{i=0}^{n}t_{i}X^{i}\right)_{n}+\left(s\cdot\sum_{i>n}t_{i}X^{i-n-1}\cdot X^{n+1}\right)_{n}=\left(s\cdot\sum_{i=0}^{n}t_{i}X^{i}\right)_{n}.$
Now, by using that $\cdot$ extends de product of $A[X;\sigma,\delta]$ and (2.10), we get
$\left(s\cdot\sum_{i=0}^{n}t_{i}X^{i}\right)_{n}=\left(\left(\sum_{i=0}^{q_{n}(t)}s_{i}X^{i}\right)\cdot\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n}+\left(\left(\sum_{i>q_{n}(t)}s_{i}X^{i}\right)\cdot\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n}\\ =\left(\left(\sum_{i=0}^{q_{n}(t)}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n}+\left(\left(\sum_{i>q_{n}(t)}s_{i}X^{i-q_{n}(t)-1}\right)\cdot\left(X^{q_{n}(t)+1}\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n}\\ =\left(\left(\sum_{i=0}^{q_{n}(t)}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n}.$
Therefore,
$(s\cdot t)_{n}=\left(\left(\sum_{i=0}^{q_{n}(t)}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right)_{n},$
for all $n$, which implies that $s\cdot t=st$.
Finally, (2.11) is easily deduced from (2.10) and (2.7). ∎
Definition 2.5.
Assume that the ring $A[\![X;\sigma,\delta]\!]$ exists according to Definition 2.3. We say that this ring structure is compatible if (2.11) holds for every $s\in A[\![X;\sigma,\delta]\!]$.
The following theorem shows that there exists a tight connection between the existence of a reasonable product for skew power series and left fractions of skew polynomials with denominators at powers of $X^{-1}$. We use the standard notion of left or right denominator set (see, e.g., [11, Ch. II]).
Theorem 2.6.
The following conditions are equivalent for a left skew derivation $(\sigma,\delta)$ on a ring $A$.
  1. (i)
    $\delta$ is locally nilpotent;
  2. (ii)
    $A[\![X]\!]$ is endowed with a reasonable ring structure containing $A[X;\sigma,\delta]$ as a subring and such that $\Sigma=\{1,X,X^{2},\dots\}$ is a left denominator set;
  3. (iii)
    $\Sigma=\{1,X,X^{2},\dots\}$ is a left denominator set of $A[X;\sigma,\delta]$.
The ring structure mentioned in (ii) is unique. Moreover, it is compatible and its product satisfies (2.10).
Proof.
(i) $\Rightarrow$ (ii). Proposition 2.4 gives the reasonable ring structure on $A[\![X]\!]$ containing $A[X;\sigma,\delta]$ as a subring and also the uniqueness statement. Also, it is compatible and (2.10) holds.
Let us check that $\Sigma$ is a left denominator set. Since the product is reasonable, we easily get from (2.7) that $\Sigma$ is a left reversible multiplicative set in the sense of [11, pag. 52]. Given $f=\sum_{i=0}^{\infty}f_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$, observe that $f-f_{0}=hX$, for a suitable $h\in A[\![X;\sigma,\delta]\!]$. Set $n$ such that $\delta^{n}(f_{0})=0$. Then
$X^{n}f=X^{n}(f-f_{0})+Xf_{0}=X^{n}hX+\sum_{k=0}^{n}N_{k}^{n}(f_{0})X^{k}=X^{n}hX+(\sum_{k=1}^{n}N_{k}^{n}(f_{0})X^{k-1})X,$
since $\delta^{n}(f_{0})=N_{0}^{n}(f_{0})$. We see that $X^{n}f=gX$ for some $g\in A[\![X;\sigma,\delta]\!]$. From this, we easily get inductively that $\Sigma$ is a left permutable set and, therefore, a left denominator set.
(ii) $\Rightarrow$ (iii). If $\Sigma$ is a left denominator for the reasonable ring structure on $A[\![X]\!]$, then, given $f\in A[X;\sigma,\delta]$ and $m\geq 0$, then there are $s\in A[\![X]\!]$ and $n\geq 0$ such that $X^{n}f=sX^{m}$. Since $X^{n}f\in A[X;\sigma,\delta]$, we have, by (2.7), that $s\in A[X;\sigma,\delta]$. Thus, $\Sigma$ is a left denominator set of $A[X;\sigma,\delta]$.
(iii) $\Rightarrow$ (i). Assume $\Sigma$ to be a left denominator set. Given $a\in A$, there exist $g\in A[X;\sigma,\delta]$ and $n\in\mathbb{N}$ such that $X^{n}a=gX$. By (2.3),
$gX=X^{n}a=\sum_{i=0}^{n}N_{i}^{n}(a)X^{i}.$
So $\delta^{n}(a)=N_{0}^{n}(a)=0$, as we desired. ∎
Corollary 2.7.
Assume $\sigma$ to be an automorphism of $A$. Then $\Sigma$ is a right denominator set of $A[X;\sigma,\delta]$ if, and only if, $\delta^{\prime}=-\delta\sigma^{-1}$ is locally nilpotent.
Proof.
We have seen before that $A[X;\sigma,\delta]=[X;\sigma^{\prime},\delta^{\prime}]A$, where $\sigma^{\prime}=\sigma^{-1}$. The corollary follows from the right handed version of the equivalence between (i) and (iii) in Theorem 2.6. ∎
Remark 2.8.
Theorem 2.6 shows that, if $\delta^{\prime}$ is locally nilpotent, then the right $A$–module of formal power series, which we may denote by $[\![X]\!]A$, is endowed with the structure of a ring containing $[X;\sigma^{\prime},\delta^{\prime}]A$ as a subring. According to the version of Definition 2.3 for right skew derivations, this ring should be denoted by $[\![X;\sigma^{\prime},\delta^{\prime}]\!]A$. However, the rings $A[\![X;\sigma,\delta]\!]$ and $[\![X;\sigma^{\prime},\delta^{\prime}]\!]A$ might not be isomorphic in the expected way when $\delta$ is locally nilpotent as well, even thought that $A[X;\sigma,\delta]=[X;\sigma^{\prime},\delta^{\prime}]A$. In particular, $\Sigma$ could be not a right denominator set for $A[\![X;\sigma,\delta]\!]$. See Example 2.11 later.
We would like to explore the possibility of extending the ring structure from $A[\![X;\sigma,\delta]\!]$ given in Theorem 2.6 to $A(\!(X)\!)$. Even thought that the obvious candidate is the ring of left fractions $\Sigma^{-1}A[\![X;\sigma,\delta]\!]$, these fractions could be not in general properly represented as elements of $A(\!(X)\!)$ (see Example 2.11 below). In order to solve this difficulty, we will investigate when $\Sigma$ is a right denominator set of $A[\![X;\sigma,\delta]\!]$.
Lemma 2.9.
Let $(\sigma,\delta)$ be a left skew derivation on $A$ such that $A[\![X;\sigma,\delta]\!]$ exits with a compatible product. The following statements hold:
  1. (i)
    $\Sigma$ is a right permutable set of $f\in A[\![X;\sigma,\delta]\!]$, if, and only if, for each $f\in A[\![X;\sigma,\delta]\!]$, we have that
    (2.13) $fX^{m}=\sigma(s)X+\delta(s),$
    for some $s\in A[\![X;\sigma,\delta]\!]$ and $m\in\mathbb{N}$.
  2. (ii)
    $\Sigma$ is a right reversible set of $A[\![X;\sigma,\delta]\!]$ if, and only if, the only solution to
    (2.14) $0=Xs$
    is the trivial one.
Proof.
(i) By (2.11) , we have that 2.13 is equivalent to $fX^{m}=Xs=\sigma(s)X+\delta(s)$. So, by induction, we have that $\Sigma$ is a right permutable set.
(ii) If $\Sigma$ is right reversible then, from $Xs=0$ we should get that $sX^{k}=0$ for some $k$. By (2.7), $s=0$. The converse is clear. ∎
If $f=\sum_{i=0}^{\infty}f_{i}X^{i}$ and $s=\sum_{i=0}^{\infty}s_{i}X^{i}$, then (2.13) is equivalent to
(2.15) $\begin{array}[]{rl}0=&\delta(s_{0})\\ 0=&\sigma(s_{0})+\delta(s_{1})\\ \vdots\\ 0=&\sigma(s_{m-2})+\delta(s_{m-1})\\ f_{0}=&\sigma(s_{m-1})+\delta(s_{m})\\ f_{1}=&\sigma(s_{m})+\delta(s_{m+1})\\ \vdots\\ f_{i}=&\sigma(s_{m+i-1})+\delta(s_{m+i})\\ \vdots\\ \end{array}$
and (2.14) is equivalent to
(2.16) $\begin{array}[]{rl}0=&\delta(s_{0})\\ 0=&\sigma(s_{0})+\delta(s_{1})\\ \vdots\\ 0=&\sigma(s_{i-1})+\delta(s_{i})\\ \vdots\\ \end{array}$
Proposition 2.10.
Let $(\sigma,\delta)$ be any left skew derivation on $A$ such that $A[\![X;\sigma,\delta]\!]$ exists with a compatible product. Set $\Sigma=\{1,X,X^{2},\dots\}$. If $\delta$ is surjective, the following statements hold:
  1. (i)
    $\Sigma$ is a right permutable set of $A[\![X;\sigma,\delta]\!]$.
  2. (ii)
    $\Sigma$ is not a right reversible set of $A[\![X;\sigma,\delta]\!]$.
Proof.
(i) Let us prove that $\Sigma$ is a right permutable set of $A[\![X;\sigma,\delta]\!]$. Given $f=\sum_{i=0}^{\infty}f_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$, we need to find $s=\sum_{i=0}^{\infty}s_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$ and $m\in\mathbb{N}$, such that satisfy (2.15), by Lemma 2.9.
Since $\delta$ is surjective, given $f$, define $s\in A[\![X;\sigma,\delta]\!]$ recursively by $\delta(s_{0})=f_{0}$ and $\delta(s_{i})=f_{i}-\sigma(s_{i-1})$ for $i\geq 1$. This series $s$ is a solution to (2.15) for $m=0$.
(ii) By the subjectivity of $\delta$, we can define $s\in A[\![X;\sigma,\delta]\!]$ recursively by $s_{0}=1$ and $s_{i+1}\in A$ such that $\delta(s_{i+1})=-\sigma(s_{i})$. Then $Xs=0$ and thus, $\Sigma$ is not a right reversible set of $A[\![X;\sigma,\delta]\!]$, by Lemma 2.9.
Example 2.11.
Let $\mathbb{F}[Y]$ be the commutative polynomial ring over a field $\mathbb{F}$ of characteristic 00. Take $\sigma=1$ and $\delta$ the usual derivation. Then $\delta$ is locally nilpotent and surjective and thus, by Theorem 2.6, $A[\![X;\sigma,\delta]\!]$ exists with a compatible product and $\Sigma$ is a left denominator set of $A[\![X;\sigma,\delta]\!]$. By Proposition 2.10, $\Sigma$ is a right permutable set but not a right reversible. Note that, if there exists a ring structure on $A(\!(X)\!)$ containing $A[\![X;\sigma,\delta]\!]$ as a subring and such that $X^{-1}X=1$, then $Xs=0$ forces $s=0$. So, no such a ring structure could exist in this example.
In this example, both $A[\![X;\sigma,\delta]\!]$ and $[\![X;\sigma^{\prime},\delta^{\prime}]\!]A$ do exist (see Remark 2.8) for this notation). However, there is no an isomorphism between them extending the identity on $A[X;\sigma,\delta]=[X;\sigma^{\prime},\delta^{\prime}]A$: Indeed, assume that $\alpha:A[\![X;\sigma,\delta]\!]\to[\![X;\sigma^{\prime},\delta^{\prime}]\!]A$ is any injective ring homomorphism such that $\alpha(X)=X$. We can choose, as in the proof of Proposition 2.10, $0\neq s\in A[\![X;\sigma,\delta]\!]$ such that $Xs=0$. Then $0=\alpha(Xs)=\alpha(X)\alpha(s)=X\alpha(s)$. This equality in $[\![X;\sigma^{\prime},\delta^{\prime}]\!]A$ implies that $\alpha(s)=0$, which is not possible.
Next, we will discuss when the Laurent series left $A$–module $A(\!(X)\!)$ becomes a ring containing $A[\![X;\sigma,\delta]\!]$ as a subring.
Lemma 2.12.
Let $A[\![X]\!]$ be endowed with a ring structure such that (2.7) holds. Set $\Sigma=\{1,X,X^{2},\dots\}$. The following statements are equivalent.
  1. (i)
    $A(\!(X)\!)$ has a ring structure such that $X^{i}$ and $X^{-i}$ are mutually inverse for any $i$, and $A[\![X]\!]$ is a subring;
  2. (ii)
    $A[\![X]\!]$ has a ring structure such that $\Sigma$ is a right denominator set.
In such a case, $A(\!(X)\!)$ is isomorphic to $A[\![X]\!]\Sigma^{-1}$ with a ring isomorphism whose restriction to $A[\![X]\!]$ is the canonical map $A[\![X]\!]\to A[\![X]\!]\Sigma^{-1}$.
Proof.
Assume (i). Given any $t=\sum_{i=i_{0}}^{\infty}t_{i}X^{i}\in A(\!(X)\!)$, with $i_{0}<0$, we use (2.7) and that any $X^{i}$ and $X^{-i}$ are mutually inverse to do the following computation.
$\begin{array}[]{rl}tX^{-i_{0}}=&\left(\sum_{i=i_{0}}^{-1}t_{i}X^{i}\right)X^{-i_{0}}+\left(\sum_{i=0}^{\infty}t_{i}X^{i}\right)X^{-i_{0}}\\ =&\sum_{i=i_{0}}^{-1}t_{i}X^{i-i_{0}}+\sum_{i=0}^{\infty}t_{i}X^{i-i_{0}}\\ =&\sum_{i=i_{0}}^{\infty}t_{i}X^{i-i_{0}}.\end{array}$
Then, every element of $A(\!(X)\!)$ is of the form $sX^{\ell}$ for some $s\in A[\![X]\!]$ and $\ell\in\mathbb{Z}$. Therefore, $A(\!(X)\!)$ is a right ring of fractions of $A[\![X]\!]$ with respect to $\Sigma$. By [11, II.1.4], $\Sigma$ is a right denominator set of $A[\![X]\!]$.
Conversely, asume (ii), which implies that there exists a ring of fractions $A[\![X]\!]\Sigma^{-1}$. The kernel of its canonical map $A[\![X]\!]\to A[\![X]\!]\Sigma^{-1}$ is
$\{s\in A[\![X]\!]:sX^{n}=0\text{ for some }n\geq 0\},$
which is zero by (2.7).
Every nonzero element of $A[\![X]\!]\Sigma^{-1}$ is of the form $(\sum_{i=j_{0}}^{\infty}s_{i}X^{i})X^{k_{0}}$, with $j_{0}\geq 0,k_{0}\leq 0$ and $s_{j_{0}}\neq 0$. Thus, by (2.7),
$\left(\sum_{j=j_{0}}^{\infty}s_{j}X^{j}\right)X^{k_{0}}=\left(\left(\sum_{j=0}^{\infty}s_{j+j_{0}}X^{j}\right)X^{j_{0}}\right)X^{k_{0}}=\left(\sum_{j=0}^{\infty}s_{j+j_{0}}X^{j}\right)X^{j_{0}+k_{0}}.$
Thus, every nonzero element of $A[\![X]\!]\Sigma^{-1}$ is written as $sX^{i_{0}}$ for $s\in A[\![X]\!]$ with $s_{0}\neq 0$ and $i_{0}\in\mathbb{Z}$. Let us check that this expression is unique: suppose that $(\sum_{i=0}^{\infty}s_{i}X^{i})X^{i_{0}}=(\sum_{j=0}^{\infty}t_{j}X^{i})X^{j_{0}}$ with $s_{0}\neq 0,t_{0}\neq 0$ and $i_{0},j_{0}\in\mathbb{Z}$. We may assume that $i_{0}\geq j_{0}$. Hence, $(\sum_{i=0}^{\infty}s_{i}X^{i})X^{i_{0}-j_{0}}=\sum_{j=0}^{\infty}t_{j}X^{i}$ in $A[\![X]\!]\Sigma^{-1}$. Since the canonical map is injective, the former equality also holds in $A[\![X]\!]$. By (2.7) this is possible only if $i_{0}=j_{0}$ and $\sum_{i=0}^{\infty}s_{i}X^{i}=\sum_{j=0}^{\infty}t_{j}X^{i}$.
The map sending $(\sum_{i=0}^{\infty}s_{i}X^{i})X^{i_{0}}$ onto $\sum_{i=i_{0}}^{\infty}s_{i}X^{i+i_{0}}$ gives an isomorphism of left $A$–modules from $A[\![X]\!]\Sigma^{-1}$ onto $A(\!(X)\!)$. We may use it to transfer the ring structure of $A[\![X]\!]\Sigma^{-1}$ to $A(\!(X)\!)$, which satisfies (i). Finally, the restriction of the inverse of this map to $A[\![X]\!]$ is clearly the canonical map from $A[\![X]\!]$ to $A[\![X]\!]\Sigma^{-1}$. ∎
Definition 2.13.
Assume that $A[\![X;\sigma,\delta]\!]$ exists with a compatible product. If $\Sigma=\{1,X,X^{2},\dots\}$ is a right denominator set then the ring of left skew Laurent formal power series is defined as the ring of fractions $A(\!(X;\sigma,\delta)\!)=A[\![X;\sigma,\delta]\!]\Sigma^{-1}$. According to Lemma 2.12, this ring structure can be understood to be defined on $A(\!(X)\!)$. This ring contains $A[\![X;\sigma,\delta]\!]$ and $A[X;\sigma,\delta]$ as subrings.
Assume $A[\![X;\sigma,\delta]\!]$ to be endowed with a reasonable and compatible ring structure for $(\sigma,\delta)$. The identity (2.11) extends to
(2.17) $X^{n}s=\sum_{k=0}^{n}N_{k}^{n}(s)X^{k},\quad(s\in A[\![X;\sigma,\delta]\!]).$
for additive endomorphisms of $A[\![X;\sigma,\delta]\!]$ defined recursively, for $0\leq i\leq n$, by
$N_{i}^{n+1}=\sigma N_{i-1}^{n}+\delta N_{i}^{n},\qquad(N_{-1}^{n}=0,N_{0}^{0}=id_{A[\![X;\sigma,\delta]\!]},N_{n+1}^{n}=0).$
When $\sigma$ is an automorphism, observe that, for all $s\in A[\![X;\sigma,\delta]\!]$, one deduces from (2.11) that
(2.18) $sX=X\sigma^{\prime}(s)+\delta^{\prime}(s),$
and, by induction,
(2.19) $sX^{n}=X\sum_{k=0}^{n-1}\sigma^{\prime}\delta^{\prime k}(s)X^{n-1-k}+\delta^{\prime n}(s),$
for all $n\geq 1$.
Proposition 2.14.
Let $(\sigma,\delta)$ be a left skew derivation on $A$ such that $A[\![X;\sigma,\delta]\!]$ is endowed with a reasonable and compatible ring structure. If $\sigma$ is an automorphism, then $\Sigma=\{1,X,X^{2},\dots\}$ is a right permutable set of $A[\![X;\sigma,\delta]\!]$ if, and only if, for every $s\in A[\![X;\sigma,\delta]\!]$ there exist $n\in\mathbb{N}$ and $t\in A[\![X;\sigma,\delta]\!]$ such that $\delta^{\prime n}(s)=Xt$.
Proof.
If $\Sigma$ is a right permutable set, then, given $s\in A[\![X;\sigma,\delta]\!]$, there are $n\in\mathbb{N},r\in A[\![X;\sigma,\delta]\!]$ such that $sX^{n}=Xr$. From (2.19) we get
$\delta^{\prime n}(s)=X\left(r-\sum_{k=0}^{n-1}\sigma^{\prime}\delta^{\prime k}(s)X^{n-1-k}\right).$
Conversely, if, given $s\in A[\![X;\sigma,\delta]\!]$, we have that $\delta^{\prime n}(s)=Xt$ for some $t\in A[\![X;\sigma,\delta]\!]$, then, by (2.19),
$sX^{n}=X\left(\sum_{k=0}^{n-1}\sigma^{\prime}\delta^{\prime k}(s)X^{n-1-k}+t\right).$
An inductive argument shows then that $\Sigma$ is a right permutable set for $A[\![X;\sigma,\delta]\!]$.
Theorem 2.15.
Let $(\sigma,\delta)$ be a left skew derivation on $A$ such that $A[\![X;\sigma,\delta]\!]$ is endowed with a reasonable and compatible ring structure. If $\sigma$ is an automorphism and $\delta^{{}^{\prime}}$ is nilpotent, then $\Sigma$ is a right denominator set in $A[\![X;\sigma,\delta]\!]$ and the product of $A(\!(X;\sigma,\delta)\!)$ satisfies that
(2.20) $X^{-1}s=\sum_{k=0}^{m}\sigma^{\prime}\delta^{\prime k}(s)X^{-1-k},\qquad(s\in A[\![X;\sigma,\delta]\!]).$
Proof.
$\Sigma$ is right permutable, by Proposition 2.14. Now, suppose that there exists $s\in A[\![X;\sigma,\delta]\!]$ such that $0=Xs$. Define $t=\sum_{i=0}^{\infty}t_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$ with $t_{i}=-\sigma(s_{i})$. Then, for every $i,j$ we deduce, by using (2.16), that $\delta^{{}^{\prime}i}(t_{i+j})=t_{j}$. Since $\delta^{{}^{\prime}}$ is nilpotent, we obtain that $t=0$ which implies that $s=0$ because $\sigma$ is an automorphism. By Lemma 2.9, $\Sigma$ is right reversible.
Finally, (2.20) is immediate from (2.19) ∎
Next example shows that the local nilpotency of $\delta$ and $\delta^{\prime}$, together with the right reversibility of $\Sigma$, is not sufficient to ensure that $\Sigma$ is a right denominator set.
Example 2.16.
Let $A=\mathbb{F}[Y,Z]$ where $\mathbb{F}$ is a field and $\mathbb{F}[Y,Z]$ is the commutative polynomial ring with coefficients in $\mathbb{F}$ over the variables $Y$ and $Z$. Take $\sigma=1$ and define $\delta(f(Y,Z))=Y\frac{\partial f}{\partial Z}$. It is clear that $\delta$ and $\delta^{\prime}=-\delta$ are both derivations which are locally nilpotent but not nilpotent. We claim that there is no $t=\sum_{i=0}^{\infty}t_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$ such that $\delta^{{}^{\prime}i}(t_{i+j})=t_{j}$ and $\delta^{\prime}(t_{0})=0$. Indeed, if that such $t$ exists, then $t_{0}=f(Y)$ for some polynomial in $\mathbb{F}[Y]$. Now, $t_{0}=\delta^{\prime}(t_{1})=-Y\frac{\partial f(Z)}{\partial Z}$, and thus $Y$ divides $t_{0}$. Inductively, $Y^{k}$ divides $t_{0}$ for all $k\geq 1$, which is not possible unless $t=0$. So, the only solution to (2.16) is the trivial one. Then, by Lemma 2.9, $\Sigma$ is right reversible set.
Let us see that $\Sigma$ is not a right permutable set of $A[\![X;\sigma,\delta]\!]$. Suppose that $\Sigma$ is a right permutable set, then, for $f=\sum_{i=0}^{\infty}Z^{i+1}X^{i}\in A[\![X;\sigma,\delta]\!]$, there are $s=\sum_{i=0}^{\infty}s_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$ and $k\geq 0$ such that $fX^{k}=Xs$. Then by (2.15), we have that
$Z=s_{k-1}+\delta(s_{k})\text{ and }Z^{2}=s_{k}+\delta(s_{k+1}).$
Since $\delta(s_{k})=Y\frac{\partial s_{k}}{\partial Z}$, we have that $s_{k-1}=Z$ and $s_{k}\in\mathbb{F}[Y]$. But this makes impossible that $Z^{2}=s_{k}+Y\frac{\partial s_{k+1}}{\partial Z}$. So, $\Sigma$ is not a right permutable set of $A[\![X;\sigma,\delta]\!]$.
We collect in the following theorem some information relevant for our purposes concerning cyclic convolutional codes.
Theorem 2.17.
Let $(\sigma,\delta)$ be a left skew derivation such that $\sigma$ is an automorphism and $\delta$ is locally nilpotent. Assume that there exists $m\geq 1$ such that $\delta^{\prime m}=0$. There is a unique ring structure on $A(\!(X)\!)$, denoted by $A(\!(X;\sigma,\delta)\!)$, containing $A[X;\sigma,\delta]$ as a subring and such that the multiplication obeys the rules
(2.21) $sX^{\ell}=\sum_{i=0}^{\infty}s_{i}X^{i+\ell},$
(2.22) $st=\sum_{n=0}^{\infty}\left[\left(\sum_{i=0}^{q_{n}}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n}X^{n},$
(2.23) $X^{-1}s=\sum_{k=0}^{m}\sigma^{\prime}\delta^{\prime k}(s)X^{-1-k},$
for all $\ell\in\mathbb{Z},s,t\in A[\![X;\sigma,\delta]\!]$ and $q_{n}\geq\max_{0\leq i\leq n}\{O_{n}(t_{i})\}$ for all $n\in\mathbb{N}$. Moreover, $A[\![X;\sigma,\delta]\!]$ is a subring of $A(\!(X;\sigma,\delta)\!)$.
Proof.
Proposition 2.4 guarantees that $A[\![X;\sigma,\delta]\!]$ exists with a reasonable and compatible ring structure. By Theorem 2.15, we get the ring $A(\!(X;\sigma,\delta)\!)$ which, as a left module, is $A(\!(X)\!)$ (see Lemma 2.12 and Definition 2.13). Moreover, it contains $A[\![X;\sigma,\delta]\!]$ as a subring. The product of this subring is unique by Proposition 2.4 and determines uniquely that of the ring of fractions $A(\!(X;\sigma,\delta)\!)=A[\![X;\sigma,\delta]\!]\Sigma^{-1}$. Thus, the product of any two Laurent power series is determined by (2.23), (2.22) and (2.21). ∎
Remark 2.18.
The skew Laurent series ring from [10, Proposition 1.2] is a particular case of 2.17. Observe that our construction does not require $A$ to be Noetherian nor any topological condition on $A$, $\sigma$ or $\delta$.
Finally, under the conditions of Theorem 2.17, we find the skew Laurent polynomials form a subring of $A(\!(X;\sigma,\delta)\!)$, as the following proposition proves.
Proposition 2.19.
Suppose that $\sigma$ is an automorphism, $\delta$ is locally nilpotent and $\delta^{\prime}$ is nilpotent, say of degree $m$ and let $s\in A[\![X;\sigma,\delta]\!]$. Then
(2.24) $X^{-n}s=\sum_{k=0}^{n(m-1)}\left(\sum_{\mathrel{\mathop{\kern 0.0ptk_{1}+\cdots+k_{n}=k}\limits_{k_{j}\leq m-1}}}\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n}}(s)\right)X^{-n-k}$
for each $n\geq 1$.
In particular, the Laurent polynomial left $A$–submodule $A[X,X^{-1}]$ becomes a subring of $A(\!(X;\sigma,\delta)\!)$ which will be denoted by $A[X,X^{-1};\sigma,\delta]$.
Proof.
We proceed by induction. If $n=1$, (2.24) is exactly (2.23). Now suppose that (2.24) holds for $n$. Then, by (2.23) and inductive hypothesis,
$\begin{array}[]{l}X^{-n-1}s\\ =X^{-1}\left(\mathrel{\mathop{\kern 0.0pt\mathrel{\mathop{\kern 0.0pt\sum}\limits^{n(m-1)}}}\limits_{k=0}}\left(\mathrel{\mathop{\kern 0.0pt\sum}\limits_{\mathrel{\mathop{\kern 0.0ptk_{1}+\cdots+k_{n}=k}\limits_{k_{j}\leq m-1}}}}\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n}}(s)\right)X^{-n-k}\right)\\ =\left(\mathrel{\mathop{\kern 0.0pt\mathrel{\mathop{\kern 0.0pt\sum}\limits^{n(m-1)}}}\limits_{k=0}}\left(\mathrel{\mathop{\kern 0.0pt\sum}\limits_{\mathrel{\mathop{\kern 0.0ptk_{1}+\cdots+k_{n}=k}\limits_{k_{j}\leq m-1}}}}X^{-1}\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n}}(s)\right)X^{-n-k}\right)\\ =\left(\mathrel{\mathop{\kern 0.0pt\mathrel{\mathop{\kern 0.0pt\sum}\limits^{n(m-1)}}}\limits_{k=0}}\mathrel{\mathop{\kern 0.0pt\sum}\limits_{\mathrel{\mathop{\kern 0.0ptk_{1}+\cdots+k_{n}=k}\limits_{k_{j}\leq m-1}}}}\left(\mathrel{\mathop{\kern 0.0pt\mathrel{\mathop{\kern 0.0pt\sum}\limits^{m-1}}}\limits_{i=0}}\sigma^{\prime}\delta^{i}(\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n}}(s)\right)X^{-1-i}X^{-n-k}\right)\\ =\left(\mathrel{\mathop{\kern 0.0pt\mathrel{\mathop{\kern 0.0pt\sum}\limits^{(n+1)(m-1)}}}\limits_{k=0}}\left(\mathrel{\mathop{\kern 0.0pt\sum}\limits_{\mathrel{\mathop{\kern 0.0ptk_{1}+\cdots+k_{n+1}=k}\limits_{k_{j}\leq m-1}}}}\left(\sigma^{\prime}\delta^{k_{1}}(\sigma^{\prime}\delta^{\prime k_{2}}\cdots\sigma^{\prime}\delta^{\prime k_{n+1}}(s))\right)\right)X^{-n-1-k}\right),\\ \end{array}$
as we required. ∎
Suppose that $\sigma$ is an automorphism and that both $\delta$ and $\delta^{\prime}$ are nilpotent. Choose $m\geq 1$ such that $\delta^{m}=\delta^{\prime m}=0$. Using (2.10) and (2.5), and taking that $q_{n}(a)=(n+1)m-1$ for any $a\in A$ and all $n\geq 0$ into account, we get
(2.25) $\left(\sum_{i=0}^{\infty}s_{i}X^{i}\right)a=\sum_{i=0}^{\infty}\left(\sum_{j=i}^{(i+1)m-1}s_{j}N_{i}^{j}(a)\right)X^{i},$
for all $a_{i}\in A$.
3. Modules
Given a right $A$–module $M$ and a ring extension $B$ of $A$, the additive group $M\otimes_{A}B$ is endowed with a right $B$–module structure determined by the rule
$(m\otimes_{A}b)b^{\prime}=m\otimes_{A}bb^{\prime}$
for any $m\in M,b,b^{\prime}\in B.$ This construction may be applied to $B=A[X;\sigma,\delta]$. There is a canonical isomorphism of additive groups
$M\otimes_{A}A[X;\sigma,\delta]\to M[X],$
defined on generators of the tensor product by
$m\otimes_{A}\sum_{i=0}^{n}s_{i}X^{i}\mapsto\sum_{i=0}^{n}ms_{i}X^{i}.$
The right $A[X;\sigma,\delta]$–module structure of $M\otimes_{A}A[X;\sigma,\delta]$ is transferred to $M[X]$. Explicitly, by (2.3), this right module structure on $M[X]$ is given by
$\left(\sum_{i=0}^{n}m_{i}X^{i}\right)\left(\sum_{j=0}^{\ell}f_{j}X^{j}\right)=\sum_{j=0}^{\ell}\sum_{i=0}^{n}\left(\sum_{k=i}^{n}m_{k}N_{i}^{k}(f_{j})\right)X^{i+j}.$
Assume $A[\![X;\sigma,\delta]\!]$ exists with a compatible ring structure. Let
$\pi:M\otimes_{A}A[\![X;\sigma,\delta]\!]\to M[\![X]\!],$
defined on generators of the tensor product by
$m\otimes_{A}\sum_{i=0}^{\infty}s_{i}X^{i}\mapsto\sum_{i=0}^{\infty}ms_{i}X^{i}.$
If $M$ is finitely presented as a right $A$–module, then the former map is an isomorphism (see, e.g. [11, Lemma I.13.2]). Thus, the right $A[\![X;\sigma,\delta]\!]$–module structure on $M\otimes_{A}A[\![X;\sigma,\delta]\!]$ is transferred to a right $A[\![X;\sigma,\delta]\!]$–module structure to $M[\![X]\!]$. So, it is useful to know how to compute $\pi^{-1}$. Given a set of generators $\{m_{1},\dots,m_{h}\}$ of the right $A$–module $M$, and $s=\sum_{i=0}^{\infty}s_{i}X^{i}\in M[\![X]\!]$, set, for every $i$, $s_{i}=\sum_{k=1}^{h}m_{k}s_{ki}$, for adequate $s_{ki}\in A$. Then
(3.1) $\pi^{-1}\left(\sum_{i=0}^{\infty}s_{i}X^{i}\right)=\sum_{k=1}^{h}m_{k}\otimes_{A}\sum_{i=0}^{\infty}s_{ki}X^{i}.$
As a consequence of (3.1) and (2.7), the right $A[\![X;\sigma,\delta]\!]$–module structure on $M[\![X]\!]$ satisfies the rule
(3.2) $\left(\sum_{i=0}^{\infty}s_{i}X^{i}\right)X^{n}=\sum_{i=0}^{\infty}s_{i}X^{i+n}.$
When $\delta$ is locally nilpotent, the right $A[\![X;\sigma,\delta]\!]$–structure of $M[\![X]\!]$ obeys a similar rule than the product of the skew formal series ring.
Proposition 3.1.
Assume that $\delta$ is locally finite. For any finitely presented right $A$–module $M$, there is a right $A[\![X;\sigma,\delta]\!]$–module structure on $M[\![X]\!]$ such that, by restriction of scalars, makes $M[X]$ an $A[X;\sigma,\delta]$–submodule of $M[\![X]\!]$. Moreover, for $s=\sum_{i=0}^{\infty}s_{i}X^{i}\in M[\![X]\!]$ and $t=\sum_{i=0}^{\infty}t_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$, for any sequence of non-negative integers $q_{n}$ such that $q_{n}\geq\max_{0\leq i\leq n}\{O_{n}(t_{i})\}$ we have
(3.3) $st=\sum_{n=0}^{\infty}\left[\left(\sum_{i=0}^{q_{n}}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n}X^{n}.$
Proof.
We make use of formula (3.1). Thus, given $t=\sum_{i=0}^{\infty}t_{i}X^{i}\in A[\![X;\sigma,\delta]\!]$, if we set, for any $n\in\mathbb{N}$, $q_{n}\geq\max_{0\leq i\leq n}\{O_{n}(t_{i})\},$ then we get
$(st)_{n}=\sum_{k=1}^{h}m_{k}\left[\left(\sum_{i=0}^{q_{n}}s_{ki}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n}\\ =\left[\left(\sum_{k=1}^{h}\sum_{i=0}^{q_{n}}m_{k}s_{ki}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n}=\left[\left(\sum_{i=0}^{q_{n}}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n},$
according to (2.10). ∎
Proposition 3.2 implies that, under its hypotheses, we can consider a well defined right $A(\!(X;\sigma,\delta)\!)$–module structure on $M(\!(X)\!)$.
Proposition 3.2.
Assume that $A[\![X;\sigma,\delta]\!]$ is endowed with a reasonable and compatible ring structure for $(\sigma,\delta)$. If $\Sigma=\{1,X,X^{2},\dots\}$ is a right denominator set for $A[\![X;\sigma,\delta]\!]$, then, for every finitely presented right $A$–module $M$, the canonical map $M[\![X;\sigma,\delta]\!]\to M[\![X;\sigma,\delta]\!]\Sigma^{-1}$ is injective. Moreover, every nonzero element of $M[\![X;\sigma,\delta]\!]\Sigma^{-1}$ is uniquely written as $sX^{i_{0}}$ for $s\in M[\![X;\sigma,\delta]\!]$ with $s_{0}\neq 0$ and $i_{0}\in\mathbb{Z}$.
Proof.
The kernel of the canonical map $M[\![X]\!]\to M[\![X]\!]\Sigma^{-1}$ consists of those power series $s$ such that $sX^{n}=0$ for some $n\in\mathbb{N}$. It follows from (3.2) that this kernel is zero. Following the steps of the proof of Proposition 2.12 one obtains a valid one for the present proposition. ∎
We finally record the version of Theorem 2.17 for finitely presented right $A$–modules.
Theorem 3.3.
Let $(\sigma,\delta)$ be a left skew derivation such that $\sigma$ is an automorphism and $\delta$ is locally nilpotent. Assume that there exists $m\geq 1$ such that $\delta^{\prime m}=0$. For every finitely presented right $A$–module $M$, the Laurent power series module $M(\!(X)\!)$ is endowed with the structure of a right $A(\!(X;\sigma,\delta)\!)$–module such that, by restriction of scalars, makes $M[\![X]\!]$ an $A[\![X;\sigma,\delta]\!]$–submodule of $M(\!(X)\!)$. This module structure is determined by the rules
$sX^{\ell}=\sum_{i=0}^{\infty}s_{i}X^{i+\ell},$
$st=\sum_{n=0}^{\infty}\left[\left(\sum_{i=0}^{q_{n}}s_{i}X^{i}\right)\left(\sum_{i=0}^{n}t_{i}X^{i}\right)\right]_{n}X^{n},$
$yX^{-1}t=\sum_{k=0}^{m}y\sigma^{\prime}\delta^{\prime k}(t)X^{-1-k},$
for all $y\in M,\ell\in\mathbb{Z},s\in M[\![X;\sigma,\delta]\!],t\in A[\![X;\sigma,\delta]\!]$ and $q_{n}\geq\max_{0\leq i\leq n}\{O_{n}(t_{i})\}$ for all $n\in\mathbb{N}$.
Proof.
It follows from Proposition 3.1, Proposition 3.2 and Corollary 2.15. ∎
4. General cyclic convolutional codes
Let us first recall the classical definition of convolutional code according to C. Roos. By $\mathbb{F}$ we denote any field. We have then the Laurent formal power series field $\mathbb{F}(\!(X)\!)$. For any $\mathbb{F}$–vector space $V$, the $\mathbb{F}$–vector space $V(\!(X)\!)$ becomes an $\mathbb{F}(\!(X)\!)$–vector space according to the rule
(4.1) $\left(\sum_{i}s_{i}X^{i}\right)\left(\sum_{j}t_{j}X^{j}\right)=\sum_{k}\left(\sum_{i+j=k}s_{i}t_{j}\right)X^{k},$
for all $\sum_{i}s_{i}X^{i}\in V(\!(X)\!)$ and $\sum_{j}t_{j}X^{j}\in\mathbb{F}(\!(X)\!)$. Analogously, $V[X]$ becomes an $\mathbb{F}[X]$–module.
Definition 4.1.
[8, Definition 1] A convolutional linear code of rate $k/n$ is a $k$–dimensional vector subspace of $\mathbb{F}^{n}(\!(X)\!)$ with a basis consisting polynomial sequences, that is, of polynomial vectors in $\mathbb{F}^{n}[X]$.
Our aim is to state a mathematically sound notion of cyclic convolutional code with respect to a left skew derivation $(\sigma,\delta)$ of a finite dimensional algebra $A$ in the spirit of Definition 4.1. To this end, we will use the Laurent formal power series rings $A(\!(X;\sigma,\delta)\!)$ and their modules discussed in the previous sections.
In what follows, $A$ will denote a finite-dimensional algebra over the field $\mathbb{F}$. Under this hypothesis, every locally nilpotent $\mathbb{F}$–linear endomorphism of $A$ becomes nilpotent. Along this section we will assume that $\sigma$ is an $\mathbb{F}$–algebra automorphism of $A$ and that both $\delta$ and $\delta^{\prime}$ are nilpotent $\mathbb{F}$–linear skew derivations. We thus fix an index $m$ such that $\delta^{m}=\delta^{\prime m}=0$.
Observe that the existence of this last $\mathbb{F}$–algebra structure on $A(\!(X)\!)$ is guaranteed by Theorem 2.17. Given a finite-dimensional right $A$–module $M$, the Laurent power series right $A$–module $M(\!(X)\!)$ is endowed with the structure of a right $A(\!(X;\sigma,\delta)\!)$–module by virtue of Theorem 3.3. In this way, by restriction of scalars, $M(\!(X)\!)$ is a right $A$–module. Precisely, it follows from (2.24) and (2.25), that, for $a\in A$, $s\in M(\!(X)\!)$.
[…]
Also, since $\mathbb{F}(\!(X)\!)$ is a subring of $A(\!(X;\sigma,\delta)\!)$, we see that $M(\!(X)\!)$ becomes a right $\mathbb{F}(\!(X)\!)$–vector space. By (3.3) and (2.24), this vector space structure is governed, as we might expect, by the rule (4.1), since $\delta,\sigma$ and $\delta^{\prime}$ are $\mathbb{F}$–linear.
Fix an $\mathbb{F}$–basis $\{a_{1},\dots,a_{r}\}$ of $A$.
Lemma 4.2.
A subset $\mathcal{D}$ of $M(\!(X)\!)$ is a right $A(\!(X;\sigma,\delta)\!)$–submodule of $M(\!(X)\!)$ if, an only if, it is both an $\mathbb{F}(\!(X)\!)$–vector subspace and a right $A$–submodule of $M(\!(X)\!)$.
Proof.
Let us just check that, if $\mathcal{D}$ is an $\mathbb{F}(\!(X)\!)$–vector subspace and a right $A$–submodule of $M(\!(X)\!)$, then it is a right $A(\!(X;\sigma,\delta)\!)$–module. Thus, pick $sX^{-n_{0}}\in\mathcal{D}$, where $s\in M[\![X]\!]$ and $n\geq 0$, and $t=\sum_{i}t_{i}X^{i}\in A(\!(X;\sigma,\delta)\!)$. For each $i$ write $t_{i}=\sum_{j=1}^{r}t_{ij}a_{j}$ for suitable $t_{ij}\in\mathbb{F}$. Thus, $t=\sum_{j=1}^{r}a_{j}\sum_{i}t_{ij}X^{i}$.
$sX^{-n_{0}}t=sX^{-n_{0}}\sum_{j=1}^{r}a_{j}\sum_{i}t_{ij}X^{i}\\ =s\sum_{j=1}^{r}\sum_{k=0}^{n_{0}(m-1)}\left(\sum_{k_{1}+\cdots+k_{n_{0}}=k}\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n_{0}}}(a_{j})\right)X^{-n_{0}-k}\sum_{i}t_{ij}X^{i}\\ =s\sum_{j=1}^{r}\sum_{k=0}^{n_{0}(m-1)}\left(\sum_{k_{1}+\cdots+k_{n_{0}}=k}\sigma^{\prime}\delta^{\prime k_{1}}\cdots\sigma^{\prime}\delta^{\prime k_{n_{0}}}(a_{j})\right)\sum_{i}t_{ij}X^{i-n_{0}-k}\in\mathcal{D},$
since $s\in\mathcal{D}$. ∎
We are in position of give a definition of general cyclic structure on a convolutional code extending that of [4]. We thus assume $\mathbb{F}^{n}$ to be endowed of a right $A$–module structure.
Definition 4.3.
A convolutional code $\mathcal{D}$ is said to be $(\sigma,\delta,A)$–cyclic if $\mathcal{D}$ is a right $A$–submodule of $\mathbb{F}^{n}(\!(X)\!)$. Equivalently, by Lemma 4.2, a $(\sigma,\delta,A)$–cyclic convolutional code is just a right $A(\!(X;\sigma,\delta)\!)$–submodule of $\mathbb{F}^{n}(\!(X)\!)$ having a set of generators in $\mathbb{F}^{n}[X]$.
Remark 4.4.
Assume $\delta=0$. Then the right $A$–module structure on $\mathbb{F}^{n}(\!(X)\!)$ given in (4.2) boils down to
$\left(\sum_{i\geq-n_{0}}^{\infty}s_{i}X^{i}\right)a=\sum_{i\geq-n_{0}}^{\infty}s_{i}\sigma^{i}(a)X^{i},$
which agrees with [4, (3.1)]. As a consequence, $(\sigma,0,A)$–cyclic convolutional codes are just the $(\sigma,A)$–CCCs discussed in [4].
When $\delta=0$, the map $\mathcal{D}\mapsto\mathcal{D}\cap\mathbb{F}^{n}[X]$ is a bijection between the set of all $(\sigma,A)$–CCCs codes of rate $k/n$ and the set of all $A[X;\sigma]$–submodules of $\mathbb{F}^{n}[X]$ that are $\mathbb{F}[X]$–direct summands of rank $k$ (see [4, Proposition 2]). To extend this correspondence to the case of $(\sigma,\delta,A)$–CCCs, we need to discuss some basic facts on convolutional codes, according to [4].
We need several lemmas. Given a subset $B\subseteq\mathbb{F}^{n}(\!(X)\!)$, the notation $B^{*}$ stands for the $\mathbb{F}(\!(X)\!)$–vector subspace of $\mathbb{F}^{n}(\!(X)\!)$ spanned by $B$.
Lemma 4.5.
([4, Lemma 1]) Let $V$ be an $\mathbb{F}(\!(X)\!)$–vector subspace of $\mathbb{F}^{n}(\!(X)\!)$. Then $V\cap\mathbb{F}^{n}[X]$ is an $\mathbb{F}[X]$–direct summand of $\mathbb{F}^{n}[X]$. Moreover, $V$ is a convolutional code if and only if $V=(V\cap\mathbb{F}^{n}[X])^{*}$.
Given a subset $B\subseteq\mathbb{F}^{n}[X]$, we denote by $\overline{B}$ the smallest $\mathbb{F}[X]$–direct summand of $\mathbb{F}^{n}[X]$ that contains $B$.
Lemma 4.6.
([4, Lemma 2]) If $B\subseteq\mathbb{F}^{n}[X]$, then $\overline{B}=B^{*}\cap\mathbb{F}^{n}[X]$. Moreover, the rank of the $\mathbb{F}[X]$–module $\mathbb{F}[X]B$ is equal to the dimension over $\mathbb{F}(\!(X)\!)$ of $B^{*}$.
The following extends [4, Lemma 4] to the current setting.
Lemma 4.7.
If $B$ is any right $A$-submodule of $\mathbb{F}^{n}(\!(X)\!)$, then $B^{*}$ is a right $A$-submodule of $\mathbb{F}^{n}(\!(X)\!)$.
Proof.
The elements of $B^{*}$ are of the form $b_{1}f_{1}+\cdots+b_{t}f_{t}$ where $f_{k}\in\mathbb{F}(\!(X)\!)$ and $b_{k}\in B$ for $1\leq k\leq t$. So, it suffices if we prove that $bfa\in B^{*}$ for every $a\in A$, $b\in B$ and $f=\sum_{i\geq i_{0}}f_{i}X^{i}\in\mathbb{F}(\!(X)\!)$ with $i_{0}\in\mathbb{Z}$. We consider two cases for $s$: the sum of all monomials with non negative degrees and the sum of all monomials with negative degrees. First, let $f=\sum_{i\geq 0}f_{i}X^{i}\in\mathbb{F}[\![X]\!].$
Then, since $\delta$ is nilpotent and by (2.25),
$bfa=\left(b\sum_{i\geq 0}f_{i}X^{i}\right)a=\left(\sum_{i\geq 0}b\left(\sum_{j=i}^{(i+1)m-1}f_{j}N^{j}_{i}(a)\right)\right)X^{i}.$
Since $A$ is finite dimensional over $\mathbb{F}$ we have our $\mathbb{F}$-basis $\{a_{1},...,a_{r}\}$. Then $\sum_{j=i}^{(i+1)m-1}N^{j}_{i}(a)f_{j}=\sum_{k=1}^{r}f_{ik}a_{k}$ for some $f_{ik}\in\mathbb{F}$ for each $i\geq 0$. Therefore
$bfa=\sum_{i\geq 0}b\left(\sum_{k=1}^{r}f_{ik}a_{k}\right)X^{i}=\sum_{k=1}^{r}ba_{k}\sum_{i\geq 0}f_{ik}X^{i}\in B^{*},$
since $B$ is a $A$-submodule.
[…]
Theorem 4.8.
The map $\mathcal{D}\mapsto\mathcal{D}\cap\mathbb{F}^{n}[X]$ is a bijection between the set of all $(\sigma,\delta,A)$–cyclic convolutional codes of rate $k/n$ an the set of all right $A[X;\sigma,\delta]$–submodules of $\mathbb{F}^{n}[X]$ that are $\mathbb{F}[X]$–direct summands of rank $k$ of $\mathbb{F}^{n}[X]$.
Proof.
A $(\sigma,\delta,A)$–cyclic convolutional code $\mathcal{D}$ is a right $A(\!(X;\sigma,\delta)\!)$–submodule of $\mathbb{F}^{n}(\!(X)\!)$. Since $A[X;\sigma,\delta]$ is a subring of $A(\!(X;\sigma,\delta)\!)$, one easily gets that $\mathcal{D}\cap\mathbb{F}^{[}X]$ is an $A[X;\sigma,\delta]$–submodule of $\mathbb{F}^{n}[X]$. By Lemma 4.5, $\mathcal{D}\cap\mathbb{F}^{n}[X]$ is an $\mathbb{F}[X]$–direct summand of $\mathbb{F}^{n}[X]$. Now, $\mathcal{D}$ has an $\mathbb{F}(\!(X)\!)$–basis $B$ whose elements belong to $\mathbb{F}^{n}[X]$. Observe that, according to Lemma 4.6,
$\mathcal{D}\cap\mathbb{F}^{n}[X]=\overline{B}=\overline{BF[X]},$
and the dimension $k$ of the $\mathbb{F}(\!(X)\!)$–vector space $B^{*}=\mathcal{D}$ equals to the rank of the $\mathbb{F}[X]$–module $BF[X]$. Finally, this rank is, precisely, the rank of $\overline{BF[X]}$.
By virtue of Lemma 4.5 and Lemma 4.6, the reciprocal of the map described above, sends a right $A[X;\sigma,\delta]$–submodule $\mathcal{C}$ of $\mathbb{F}^{n}[X]$ that is an $\mathbb{F}[X]$–direct summand onto $\mathcal{C}^{*}$. Note that $\mathcal{C}^{*}$ is a right $A$–submodule of $\mathbb{F}^{n}(\!(X)\!)$ by Lemma 4.7. ∎
Recall from [6] that a right ideal code is defined as a right ideal of $A[X;\sigma,\delta]$ that it is an $\mathbb{F}[X]$–direct summand. Consider the regular right structure on $A$ and any $\mathbb{F}$–vector space isomorphism $A\cong\mathbb{F}^{n}$, with $n$ the dimension of $A$ over $\mathbb{F}$. Then $\mathbb{F}^{n}$ becomes a right $A$–module isomorphic to $A$.
Corollary 4.9.
The map $\mathcal{D}\mapsto\mathcal{D}\cap A[X;\sigma,\delta]$ is a bijection between the set of all $(\sigma,\delta,A)$–cyclic convolutional codes in $A(\!(X;\sigma,\delta)\!)$ and the set of all right ideal codes.
Proof.
In view of Lemma 4.2, we may apply Theorem 4.8 to get the bijection described in the statement of the Corollary. ∎
Let us describe some concrete examples where the former cyclic convolutional codes can be built.
Example 4.10.
Let $\mathbb{F}_{4}$ be the field with eight elements described as $\mathbb{F}_{4}=\mathbb{F}_{2}(a)$, where $a^{2}+a+1=0$. Let $\tau$ be the Frobenius automorphism on $\mathbb{F}_{8}$, that is, $\tau(c)=c^{2}$ for every $c\in\mathbb{F}_{4}$. Set $A=\mathcal{M}_{2}(\mathbb{F}_{4})$, the ring of $2\times 2$ matrices over $\mathbb{F}_{4}$, and consider the $\mathbb{F}_{2}$-automorphism $\sigma:A\to A$ defined as the component-by-component extension of $\tau$ to $A$. That is, $\sigma$ is given by
(4.3) $\sigma\left(\begin{matrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{matrix}\right)=\left(\begin{matrix}\tau(x_{0})&\tau(x_{1})\\ \tau(x_{2})&\tau(x_{3})\end{matrix}\right)=\left(\begin{matrix}x_{0}^{2}&x_{1}^{2}\\ x_{2}^{2}&x_{3}^{2}\end{matrix}\right)\text{ for every }\left(\begin{matrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{matrix}\right)\in A.Indeed,$
We can also set the inner $\sigma$-derivation $\delta:A\rightarrow A$ given by $\delta(X)=MX-\sigma(X)M$ for $X\in A$, where
$M=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$
For any $x_{0},x_{1},x_{2},x_{3}\in\mathbb{F}_{4}$,
$\begin{split}\delta\begin{pmatrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{pmatrix}&=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\begin{pmatrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{pmatrix}+\begin{pmatrix}x_{0}^{2}&x_{1}^{2}\\ x_{2}^{2}&x_{3}^{2}\end{pmatrix}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\\ &=\begin{pmatrix}x_{2}&x_{3}\\ 0&0\end{pmatrix}+\begin{pmatrix}0&x_{0}^{2}\\ 0&x_{2}^{2}\end{pmatrix}\\ &=\begin{pmatrix}x_{2}&x_{0}^{2}+x_{3}\\ 0&x_{2}^{2}\end{pmatrix}.\end{split}$
$\begin{split}\delta^{2}\begin{pmatrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{pmatrix}&=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\begin{pmatrix}x_{2}&x_{0}^{2}+x_{3}\\ 0&x_{2}^{2}\end{pmatrix}+\begin{pmatrix}x_{2}^{2}&x_{0}+x_{3}^{2}\\ 0&x_{2}\end{pmatrix}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\\ &=\begin{pmatrix}0&x_{2}^{2}\\ 0&0\end{pmatrix}+\begin{pmatrix}0&x_{2}^{2}\\ 0&0\end{pmatrix}\\ &=\begin{pmatrix}0&2x_{2}^{2}\\ 0&0\end{pmatrix}=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}.\end{split}$
Then $\delta$ is nilpotent. Notice that $\sigma(M)=M$, so $\delta\sigma=\sigma\delta$ and thus, $\delta^{\prime}$ is nilpotent. Hence $A(\!(X;\sigma,\delta)\!)$ exists.
Example 4.11.
[6, Example 5.1] Let $\mathbb{F}_{4}=\mathbb{F}_{2}(a)$ and $A=\mathbb{F}G$ the group algebra of a cyclic group $G$ of order $5$. Define $\sigma:G\to G$ via $\sigma(a)=a^{2}$, which is a group automorphism on $G$. By extending $\sigma$ linearly over $\mathbb{F}$, $\sigma$ gives an $\mathbb{F}$–algebra automorphism on $A$. Now, let $\delta$ be the inner $\sigma$-derivation induced by $1$, i.e., $\delta(a)=a+\sigma(a)$. Let us check that $\delta$ is nilpotent. Indeed, for any $a\in G$,
$\begin{array}[]{rcl}\delta(a)&=&a+a^{2}\\ \delta^{2}(a)&=&a+a^{2}+a^{2}+a^{2}=a+a^{4}\\ \delta^{3}(a)&=&a+a^{4}+a^{2}+a^{8}=a+a^{4}+a^{2}+a^{3}\\ \delta^{4}(a)&=&a+a^{2}+a^{3}+a^{4}+a^{2}+a^{4}+a^{6}+a^{8}=0.\\ \end{array}$
Since $\sigma(1)=1$, one gets that $\sigma\delta=\delta\sigma$ and, thus, $\delta^{\prime}$ is also nilpotent. Therefore, $A(\!(X;\sigma,\delta)\!)$ exists, and the codes described in [6, Example 5.1] are $(\sigma,\delta,A)$–CCCs, by Theorem 4.8.
The following example shows that, even when working with the same ring, the same automorphism $\sigma$, and an inner $\sigma$-derivation, it does not follow that $\delta$ and $\delta^{\prime}$ are nilpotent.
Example 4.12.
Consider the assumptions of Example 4.10. Now take
$M=\begin{pmatrix}0&0\\ 0&a\end{pmatrix},$
then
$\begin{split}\delta\begin{pmatrix}0&0\\ 0&a\end{pmatrix}&=\begin{pmatrix}0&0\\ 0&a\end{pmatrix}\begin{pmatrix}0&0\\ 0&a\end{pmatrix}+\begin{pmatrix}0&0\\ 0&a^{2}\end{pmatrix}\begin{pmatrix}0&0\\ 0&a\end{pmatrix}\\ &=\begin{pmatrix}0&0\\ 0&a^{2}\end{pmatrix}+\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\\ &=\begin{pmatrix}0&0\\ 0&a\end{pmatrix}.\end{split}$
and, since $\sigma^{-1}=\sigma$, we have that
$\begin{split}\delta^{\prime}\begin{pmatrix}0&0\\ 0&a\end{pmatrix}&=\begin{pmatrix}0&0\\ 0&a\end{pmatrix}.\end{split}$
Then neither $\delta$ nor $\delta^{\prime}$ is nilpotent.
Our final example shows that the nilpotence of $\delta$ needs not imply that of $\delta^{\prime}$.
Example 4.13.
Let $A=\mathbb{F}[Y,Z]/I$ where $\mathbb{F}$ is a field and $I$ is the ideal of the commutative polynomial ring $\mathbb{F}[Y,Z]$ generated by $\{Y^{2},Z^{2},YZ\}$. Set $y=Y+I,z=Z+I$ and $\sigma$ the $\mathbb{F}$–algebra automorphism of $A$ defined by $\sigma(1)=1,\sigma(y)=y+z$ and $\sigma(z)=z$. Let us define the $\mathbb{F}$–linear map $\delta$ as $\delta(1)=0,\delta(y)=0$ and $\delta(z)=y$. Let us verify that $\delta$ is a $\sigma$-derivation:
$\begin{array}[]{l}\sigma(y)\delta(y)+\delta(y)y=(y+z)0+0y=0=\delta(y^{2});\\ \sigma(z)\delta(z)+\delta(z)z=zy+yz=0=\delta(z^{2});\\ \sigma(y)\delta(z)+\delta(y)z=(y+z)y+0y=0=\delta(yz);\\ \sigma(z)\delta(y)+\delta(z)y=y0+yy=0=\delta(zy),\\ \end{array}$
as we required. Clearly, $\delta^{2}=0$.
Now,
$\delta^{\prime}(y)=-\delta(\sigma^{-1}(y))=-\delta(y-z)=y.$
Hence $\delta^{\prime}$ is not nilpotent.